A scooter company gives the following specifications about its product.
Weight of the scooter - 95 kg
Maximum speed – 60 km/h
Maximum engine power – 3.5 hp
Pick up time to get the maximum speed – 5 s
Check the validity of these specifications.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. The specification given by the company are U = 0, m = 95 kg, Pm = 3.5 hp Vm = 60 km/h = 50/3 m/sec t base m = 5 sec So, the maximum acceleration that can be produced is given by, a = (50/3) – 0/5 = 10/3 So, the driving force is given by F = ma = 95 * 10/3 = 950/3 N So, the velocity that can be attained by maximum h.p. white supplying 950/3 will be V = p/F ⇒ v = 3.5 * 746 * 5/950 = 8.2 m/sec. Because, the scooter can reach a maximum of 8.s m/sec while producing a force of 950/3 N, the specifications given are some what over claimed.